OpenStudy (anonymous):

Write the equation of the line that passes through (–3, 4) with a slope of 2 in point-slope form. y – 4 = 2(x + 3) y + 4 = 2(x – 3) y + 3 = 2(x – 4) y – 3 = 2(x + 4)

OpenStudy (anonymous):

point-slope form is y-y1=m(x-x1)

OpenStudy (anonymous):

@Luigi0210 @Compassionate guys :)

OpenStudy (compassionate):

I am investigating the disappearance of all the cats in the city... my living room is full of cats... that means...

OpenStudy (anonymous):

xD ?

OpenStudy (luigi0210):

Luigi got kidnapped

OpenStudy (anonymous):

please helpppp :)

OpenStudy (luigi0210):

@heather040200 pretty much did our work >.>

OpenStudy (compassionate):

You want to use the equation y - y1 = m(x - x1) Ya' got that foxy mama?

OpenStudy (anonymous):

i dont know what to do after that guys i told you i'm kinda dumb when it comes to this hahahaha

OpenStudy (compassionate):

Ehh, e-everyone stay calm... because we're all doomed!

OpenStudy (luigi0210):

No you're not dumb..

OpenStudy (luigi0210):

D: @Compassionate what did you do now? -_-

OpenStudy (compassionate):

Your points are: (-3, 4) and your slope is 2 y1 = 4 x1 = -3 m = 2 now plug it into y - y1 = m(x - x1)

OpenStudy (anonymous):

so the second one would be your answer

OpenStudy (luigi0210):

We gotta give Heather credit for that.

OpenStudy (anonymous):

i knew that thanks heather

OpenStudy (anonymous):

No problem :)

OpenStudy (anonymous):

i have another one thats similar, wanna try it out while we're all here? it's my last one :)

OpenStudy (luigi0210):

@heather040200 you might wanna double check your choice.. unless I'm completely losing my math skills

OpenStudy (compassionate):

Hoohah! Lets do this foxy mama! Now remember, I do my best work when I'm being worshipped as a god.

OpenStudy (anonymous):

crap @Luigi0210 hahaahhahahaha

OpenStudy (luigi0210):

I get stupider everyday >.>

OpenStudy (anonymous):

im gonna go with heather on this one, sorry Luigi xD you can redeem yourself on the next one

OpenStudy (anonymous):

Write the equation of the line that passes through (1, 3) and (4, 4) in standard form. 3x – 3y = –8 x – 3y = –9 x – 3y = –8 3x + y = 9

OpenStudy (anonymous):

ya i was gonna say with what compassion gave i would say the second one

OpenStudy (luigi0210):

>.<

OpenStudy (anonymous):

that's my last one guys let's finish strong :)

OpenStudy (anonymous):

ok i believe for this one you need to find the slope first slope is \[\frac{ y2-y1 }{ x2-x1 }\]

OpenStudy (luigi0210):

Luigi's always ignored >.>

OpenStudy (anonymous):

alright heather

OpenStudy (compassionate):

Hey, Sweet jellybean. You smell kinda pretty, wanna smell me? Hoohah! Use the equation Heather gave you. That's how you find the slope! Then after you have the slope, plug it into the equation: y - y1 = m(x - x1)

OpenStudy (anonymous):

agree with compassionate :)

OpenStudy (anonymous):

okay i plugged it in

OpenStudy (compassionate):

Okay, can you show me your work sweet-honey.

OpenStudy (anonymous):

@Compassionate you keep changing your picture haah

OpenStudy (anonymous):

too much work :P

OpenStudy (anonymous):

well at least tell what you got LOL

OpenStudy (anonymous):

i lied ahahahhaa

OpenStudy (compassionate):

I keep finding one that suits my needs! Now that you know I'm pretty, and you're pretty, why don't you say we go back to my place at each other?

OpenStudy (compassionate):

Show me the slope you gotta pretty blackmamba mama.

OpenStudy (anonymous):

you're a girl?

OpenStudy (compassionate):

Of course.

OpenStudy (anonymous):

i thought you're name was stephen

OpenStudy (compassionate):

Go re-read that again there little lady.

OpenStudy (anonymous):

oh alright

OpenStudy (anonymous):

but im working out my problem so hold on

OpenStudy (compassionate):

You take your sweet time hotty-mitotty.

OpenStudy (anonymous):

so the slope is 1/3

OpenStudy (anonymous):

i think?

OpenStudy (anonymous):

so is it y - 3 = 1/3 (x-1) ????

OpenStudy (anonymous):

idk guys

OpenStudy (compassionate):

So if the slope is 1/3 That means: (–3, 4) y1 = 4 x1 = -3 m = 1/3 Now plug it into y - y1 = m(x - x1)

OpenStudy (luigi0210):

Haha, hello there

OpenStudy (uri):

@Compassionate You're a girl since when,LIAR :O

OpenStudy (compassionate):

Luigi is back from the grave!

OpenStudy (compassionate):

Ha ha, I've always been a girl. Ask Preetha and Boston. Look at my ID. They've seen it.

OpenStudy (uri):

@Compassionate You told me you're a guy..:(

OpenStudy (compassionate):

I also told you that me and my brother use this account. :)

OpenStudy (compassionate):

Stop talking to my bro >.>

OpenStudy (anonymous):

ok i plugged it in

OpenStudy (anonymous):

y-4=1/3(x-(-3)

OpenStudy (uri):

No you never told me that -______- @Compassionate Stop trolling,We only talk about sketches.

OpenStudy (compassionate):

Huh? Sketches? .-. What are you talking about?

OpenStudy (anonymous):

so the answer is A?

OpenStudy (uri):

@Compassionate Go sleep :) Ttyl.

OpenStudy (compassionate):

Good job, Pam! :)

OpenStudy (anonymous):

yay finally.

OpenStudy (anonymous):

wait guys!!! one more omg

OpenStudy (compassionate):

Now- I can view over your work really fast. I am taking your word that you got it all correct and you're solving off of your own logic. If you want, I will go back over your work and check.

OpenStudy (anonymous):

yes please.

OpenStudy (compassionate):

You're finding the slope through the slope formula of (y2-y1)/(x2-x1) but the equation of a line can be found by using the point-slope formula, given by y-y0 = m(x-x0) you already know that m = 1/3 so you have part of the equation, the other part is just one of the points that it passes through, either one will do, but I'll do both just to show you that you end up with the same equation. So for the first one (1, 3) y - 4 = 1/3(x - 4) so now we simplify y - 4 = (1/3)x - 4/3 y = (1/3)x - 4/3 + 4 y = (1/3)x + 2-2/3 so the equation of your line is the above result, your slope is 1/3 which you already know, and your y-intercept is 2 & 2/3 or 8/3. If we were to multiply everything by 3 removing the denominator, your equation would be 3y = x + 8 So now if we use the other point (1,3), the initial equation would look like this y - 3 = 1/3(x - 1) so now simplifying y - 3 = (1/3)x - 1/3 y = (1/3)x -1/3 + 3 y = (1/3)x + 2-2/3 which is what we had before with the other point, so now you can see that the point-slope formula will work for any point along the line, and give you the same equation, just as long as you know the slope. Hope that helps.

OpenStudy (anonymous):

so was i right?

OpenStudy (compassionate):

Yes, yes you were! (Ain't Yahoo.Answers lovely!?)

OpenStudy (anonymous):

hahahaha! thank you i have one more guys

OpenStudy (compassionate):

Anything for a pretty mama like yourself!

OpenStudy (anonymous):

Write the equation of the line that passes through (1, 3) and (4, 4) in standard form. 3x – 3y = –8 x – 3y = –9 x – 3y = –8 3x + y = 9

OpenStudy (compassionate):

This is the same exact question you just posted :)

OpenStudy (anonymous):

oh! hhahahahaah

OpenStudy (luigi0210):

I love you guys >.>

OpenStudy (anonymous):

wait so that one is A?

OpenStudy (anonymous):

i love you too luigi @Luigi0210

OpenStudy (compassionate):

Ha ha, Pam, you're just adorable. I'm dying over here. You just got the answer to this one not even 2 minutes ago.

OpenStudy (anonymous):

Im just confused because the whole time i thought we were working out this question: Write the equation of the line that passes through (–3, 4) with a slope of 2 in point-slope form. y – 4 = 2(x + 3) y + 4 = 2(x – 3) y + 3 = 2(x – 4) y – 3 = 2(x + 4)

OpenStudy (luigi0210):

xD I feel the love

OpenStudy (anonymous):

hahahahahha

OpenStudy (anonymous):

oh its b nevermind :)

OpenStudy (compassionate):

Ha ha, we just worked that out XD This is too great. Someone promote Pam to administrator.

OpenStudy (anonymous):

hahahhaa sorry! but i got a 50 on my assignment so im doing it again